# M2IST: Multi-Modal Interactive Side-Tuning for Memory-efficient Referring Expression Comprehension

### v1

@@ Abstract
Referring expression comprehension (REC) is a vision-language task to locate a target object in an image based on a language expression. Fully fine-tuning general-purpose pre-trained models for REC yields impressive performance but becomes increasingly costly. Parameter-efficient transfer learning (PETL) methods have shown strong performance with fewer tunable parameters. However, applying PETL to REC faces two challenges: (1) insufficient interaction between pre-trained vision and language encoders, and (2) high GPU memory usage due to gradients passing through both heavy encoders. To address these issues, we present M2IST: Multi-Modal Interactive Side-Tuning with M3ISAs: Mixture of Multi-Modal Interactive Side-Adapters. During fine-tuning, we keep the pre-trained vision and language encoders fixed and update M3ISAs on side networks to establish connections between them, thereby achieving parameter- and memory-efficient tuning for REC. Empirical results on three benchmarks show M2IST achieves the best performance-parameter-memory trade-off compared to full fine-tuning and other PETL methods, with only 3.14M tunable parameters (2.11% of full fine-tuning) and 15.44GB GPU memory usage (39.61% of full fine-tuning). Source code will soon be publicly available.
M2IST: Multi-Modal Interactive Side-Tuning for Memory-efficient Referring Expression Comprehension
Xuyang Liu1thanks: Equal contribution. Corresponding author.  Ting Liu2∗  Siteng Huang3 Yue Hu2  Quanjun Yin2  Donglin Wang3  Honggang Chen1† 1Sichuan University  2National University of Defense Technology  3Westlake University liuxuyang@stu.scu.edu.cn  {liuting20,huyue11}@nudt.edu.cn  yin_quanjun@163.com {huangsiteng,wangdonglin}@westlake.edu.cn  honggang_chen@scu.edu.cn
@@
Figure 1: Comparison of (a) fully fine-tuning, (b) Adapter-tuning, and (c) our M2IST for REC. By updating 3.19M encoder parameters (2.11% of (a)) and requiring 15.44GB of GPU memory (39.61% of (a)), M2IST achieves comparable or even superior performance compared to fully fine-tuning (e.g., RefCOCO val Yu et al. (2016)).
1 Introduction
@@
Referring expression comprehension (REC) is one of the most challenging vision-language tasks, aiming to locate a specific object in an image based on a given referring expression Yu et al. (2018); Yang et al. (2019); Deng et al. (2021); Zhu et al. (2022); Wu et al. (2023). Recent studies Deng et al. (2021); Sun et al. (2022); Huang and Satoh (2023); Kim et al. (2024) have shown impressive performance by fine-tuning general-purpose pre-trained models for the task. However, fully fine-tuning these pre-trained models is computationally expensive when adapting to a new REC dataset (see Figure 1 (a)). Additionally, fine-tuning on limited REC data can lead to catastrophic forgetting and overfitting.
@@
Recently, parameter-efficient transfer learning (PETL) methods Houlsby et al. (2019); Hu et al. (2022); Jia et al. (2022) have been proposed to address similar issues by updating only a small set of parameters to efficiently adapt pre-trained models to downstream tasks. Adapter-tuning Houlsby et al. (2019), a typical PETL method, has achieved great success across diverse downstream tasks Yuan et al. (2023); Cao et al. (2024). It typically inserts a tunable lightweight bottleneck-shaped module sequentially into each frozen backbone layer. Most transformer-based REC models Deng et al. (2021); Sun et al. (2022); Zhang et al. (2023) use pre-trained Vision Encoder and Language Encoder to separately extract image and text features, which are then integrated to form multi-modality features for reasoning. A straightforward approach to apply adapter-tuning for REC is to insert the adapters into the transformer encoder layers to enhance fine-tuning efficiency (see Figure 1 (b)). However, this introduces two significant challenges: (1) Updating inserted adapters still requires backpropagation through the large pre-trained encoders, placing a heavy burden on GPU memory (see Figure 1 (b)). (2) The Vision and Language Encoders, pre-trained separately with different structures and data, lack cross-modality interaction in their shallow layers when vanilla adapters are inserted, leading to sub-optimal vision-language alignment. This issue is especially problematic for predicting referred objects with complex semantics, such as human actions and spatial relations.
@@
To address these challenges, we propose a novel Multi-Modal Interactive Side-Tuning (M2IST) method that effectively strengthens vision-language alignment and enables parameter- and memory-efficient transfer to REC within the unified interactive side networks (see Figure 1 (c)). Specifically, we introduce Mixture of Multi-Modal Interactive Side-Adapters (M3ISAs), which incorporate Vision Expert Adapters (VEA), Language Expert Adapters (LEA), and Interaction Expert Adapters (IEA) into the side networks in parallel with the heavy encoders. VEA and LEA transfer pre-trained single-modality knowledge to the REC domain. IEA utilizes a linear layer for weight-sharing between image and text features, enabling progressive interaction between the referring sentence and input image. This interaction aggregates multi-grained information from different modalities at shallow layers of the model, facilitating deep multi-modal fusion in deeper layers for improved reasoning. This elegant design achieves parameter- and memory-efficient intra- and inter-modality representation transfer for REC.
@@
We conduct extensive experiments on RefCOCO Yu et al. (2016), RefCOCO+ Yu et al. (2016), and RefCOCOg Mao et al. (2016); Nagaraja et al. (2016) to demonstrate the effectiveness and efficiency of M2IST for REC. Experimental results show that M2IST achieves the optimal performance-parameter-memory trade-off compared to most full fine-tuning methods and other PETL methods. Following our M2IST method, a standard transformer-based REC model can reduce 97.89% tunable encoder parameters and only require 39.61% of the GPU memory needed for full fine-tuning, while achieving competitive performance (see Figure 1). With the sufficient vision-language interaction strengthened by our M3ISAs, our method can accurately locate the referred objects for various complex cases, such as human actions and spatial relations (see Figure 4).
2 Related Work
2.1 Referring Expression Comprehension
Referring expression comprehension (REC) Yu et al. (2018); Deng et al. (2021); Zhu et al. (2022); Han et al. (2024b) aims to locate specific objects in images based on textual descriptions. Early methods Yu et al. (2018); Liu et al. (2019); Chen et al. (2019) follow a two-stage pipeline that first uses a pre-trained object detector Ren et al. (2015) to generate a set of sparse object proposals, which are then ranked by their similarity to the textual description. However, these two-stage methods heavily rely on the quality of the object proposals and cannot directly predict the referred object region. Recently, one-stage anchor-based methods Yang et al. (2019); Liao et al. (2020a); Yang et al. (2020); Ye et al. (2021) have been introduced to eliminate the proposal generation step, directly predicting the object bounding box from the pre-defined dense anchors. More recently, transformer-based methods Deng et al. (2021); Du et al. (2022); Zhu et al. (2022); Sun et al. (2022); Zhang et al. (2023) have shown superior performance by implicitly modeling cross-modality relationships in a unified architecture. As REC models continue to scale up, their performance has improved. However, this performance gain comes at the cost of increased computational cost, demanding larger GPU memory for parameter fitting (see Figure 1 (a)).
2.2 Parameter-efficient Transfer Learning
Parameter-efficient transfer learning (PETL) Houlsby et al. (2019); Hu et al. (2022); Jia et al. (2022); Chen et al. (2022); Han et al. (2024a) has emerged as a promising alternative to fully fine-tuning pre-trained models for downstream tasks. By updating only a minimal subset of parameters, PETL methods balance performance and computational efficiency. Recent PETL methods can be classified into two types: (1) Updating additional parameters in modules inserted into the model (i.e., Adapters) Houlsby et al. (2019); Chen et al. (2022); Liu et al. (2024b) or appended to the input data Jia et al. (2022); Huang et al. (2023); Xin et al. (2024); (2) Decomposing weight matrices into two low-rank matrices and updating only the small factorization matrices (e.g., LoRA) Hu et al. (2022). There is increasing interest in adapter-based PETL methods for vision-language tasks Jiang et al. (2022); Xu et al. (2023); Yuan et al. (2023); Wang et al. (2024); Liu et al. (2024a); Cao et al. (2024), which aim to achieve effective cross-modality interaction while maintaining parameter efficiency. However, existing PETL methods still face substantial GPU memory consumption during the fine-tuning stage, as gradients must propagate through the heavy pre-trained encoders for REC (see Figure 1 (b)).
2.3 Memory-efficient Transfer Learning
Memory-efficient transfer learning (METL) Sung et al. (2022); Fu et al. (2024); Zhang et al. (2024) aims to reduce memory costs on GPUs during fine-tuning. Existing METL methods typically employ a side network for single-modality knowledge transfer, focusing on either NLP Sung et al. (2022); Zhang et al. (2024) or CV Fu et al. (2024); Tang et al. (2024) downstream tasks. However, these METL methods lack sufficient cross-modality interaction between vision and language representations, which is crucial for REC. In this work, our M2IST bridges the pre-trained vision and language encoders in unified interactive side networks, facilitating parameter- and memory-efficient transfer to the REC task (see Figure 1 (c)).
3 Methodology
3.1 Base Architecture
@@
Figure 2: Overall architecture of M2IST. M2IST freezes the pre-trained Vision Encoder (blue branch) and Language Encoder (green branch), while updating M3ISAs on side networks (pink branch). M3ISAs comprise IEA for bridging the pre-trained dual encoders to enable cross-modality interactions, and VEA/LEA for transferring pre-trained single-modality representations to adapt to the REC domain. By avoiding backpropagation through the heavy encoders (red dashed arrow), M2IST enables parameter- and memory-efficient tuning for REC.
@@
We apply a standard transformer-based REC model as our base architecture, shown in Figure 1 (a), which comprises: (1) a Vision Encoder, (2) a Language Encoder, and (3) a Vision-language Encoder. Our training objective follows most transformer-based REC methods and is detailed in Appendix C.
@@
Vision Encoder. We adopt a DETR-based Carion et al. (2020) encoder as our Vision Encoder, which comprises a ResNet He et al. (2016) and a stack of transformer encoder layers to encode the image into high-quality vision embeddings. Specifically, given an input image $\bm{z}_{0}\in\mathbb{R}^{H_{0}\times W_{0}\times 3}$, the ResNet is utilized to generate a 2D feature map $\bm{z}\in\mathbb{R}^{H\times W\times C}$, where $H_{0}$ and $W_{0}$ denote the height and width of the input image, $H=\frac{H_{0}}{32}$, $W=\frac{W_{0}}{32}$, and $C=2048$ represents the channel dimension. Then, a $1\times 1$ convolutional layer reduces the $C$ to $C_{v}=256$, producing $\bm{z}^{\prime}\in\mathbb{R}^{H\times W\times C_{v}}$. We flatten the feature map $\bm{z}^{\prime}$ into a sequence of 1D vectors (i.e., vision tokens) $\bm{z}_{v}\in\mathbb{R}^{N_{v}\times C_{v}}$, where $N_{v}=H\times W$ indicates the number of tokens. Sequentially, these vision tokens added with positional encodings are fed into a stack of 6 transformer encoder layers, which then output the enhanced vision embeddings $\bm{f}_{v}\in\mathbb{R}^{N_{v}\times C_{v}}$ incorporating global context of the image.
@@
Language Encoder. We employ an off-the-shelf language model BERT Devlin et al. (2018), comprising a stack of transformer encoder layers, as our Language Encoder. Specifically, given the input text, each word ID is converted into a one-hot vector, which is then tokenized into a sequence of language tokens. These language tokens, concatenated with a [CLS] token at the beginning and a [SEP] token at the end, are input to 12 transformer encoder layers to sequentially model contextual relationships. Similar to the Vision Encoder, Language Encoder finally outputs the enhanced language embeddings $\bm{f}_{l}\in\mathbb{R}^{N_{l}\times C_{l}}$, where $N_{l}$ and $C_{l}=768$ represent the number and channel dimension of language tokens, respectively.
@@
Vision-language Encoder. We use a transformer-based encoder Vaswani et al. (2017) as our Vision-language Encoder (V-L Encoder) to thoroughly fuse the multi-modality embeddings and predict the bounding box of the referred object. Specifically, the enhanced vision embeddings $\bm{f}_{v}\in\mathbb{R}^{N_{v}\times C_{v}}$ and language embeddings $\bm{f}_{l}\in\mathbb{R}^{N_{l}\times C_{l}}$ are first projected into the joint embeddings $\bm{f^{\prime}}_{v}\in\mathbb{R}^{N_{v}\times C_{p}}$ and $\bm{f^{\prime}}_{l}\in\mathbb{R}^{N_{l}\times C_{p}}$, sharing the same channel dimension $C_{p}=256$. The joint embeddings, along with a learnable [REG] token, are then fed into a stack of 6 transformer encoder layers to fuse the cross-modality embeddings and output the [REG] token. Finally, a prediction head, implemented as a Multi-layer Perceptron with two 256-dim hidden layers and a linear output layer, receives the [REG] token and regresses it to the 4-dim box coordinates for the referred object.
3.2 Multi-Modal Interactive Side-Tuning

### v2

# M2IST: Multi-Modal Interactive Side-Tuning for Efficient Referring Expression Comprehension

@@ Abstract
Referring expression comprehension (REC) is a vision-language task to locate a target object in an image based on a language expression. Fully fine-tuning general-purpose pre-trained vision-language foundation models for REC yields impressive performance but becomes increasingly costly. Parameter-efficient transfer learning (PETL) methods have shown strong performance with fewer tunable parameters. However, directly applying PETL to REC faces two challenges: (1) insufficient multi-modal interaction between pre-trained vision-language foundation models, and (2) high GPU memory usage due to gradients passing through the heavy vision-language foundation models. To this end, we present M2IST: Multi-Modal Interactive Side-Tuning with M3ISAs: Mixture of Multi-Modal Interactive Side-Adapters. During fine-tuning, we keep the pre-trained uni-modal encoders fixed, updating M3ISAs on side networks to progressively connect them, enabling more comprehensive vision-language alignment and efficient tuning for REC. Empirical results reveal that M2IST achieves an optimal balance between performance and efficiency compared to most full fine-tuning and other PETL methods. With M2IST, standard transformer-based REC methods present competitive or even superior performance compared to full fine-tuning, while utilizing only 2.11% of the tunable parameters, 39.61% of the GPU memory, and 63.46% of the fine-tuning time required for full fine-tuning.
Index Terms:
Vision-language foundation models, parameter-efficient transfer learning, referring expression comprehension.
@@
Figure 1: Comparison of (a) fully fine-tuning, (b) Adapter-tuning, and (c) our M2IST for REC. By only using 2.11% of the tunable parameters, 39.61% of the GPU memory, and 63.46% of the fine-tuning time, M2IST achieves comparable or even superior performance compared to fully fine-tuning.
I Introduction
@@
Referring expression comprehension (REC) is one of the most challenging vision-language tasks, aiming to locate a specific object in an image based on a given referring expression [1, 2, 3, 4, 5, 6]. Recent studies [3, 7, 8, 9] have demonstrated impressive performance by fine-tuning general-purpose vision-language foundation models for this task. However, fully fine-tuning these pre-trained models is computationally expensive when adapting to a new dataset (see Figure 1 (a)). Additionally, fine-tuning on limited REC data may lead to catastrophic forgetting and overfitting, as recently evaluated in other vision-language tasks [10, 11].
@@
Recently, parameter-efficient transfer learning (PETL) methods [12, 13, 14, 15] have been proposed to address similar issues by updating only a small set of parameters to efficiently adapt pre-trained models to downstream tasks. Adapter-tuning [12], a prominent PETL method, has demonstrated significant success across various vision-language foundation models as well as downstream tasks [11, 16, 17]. It typically inserts a tunable lightweight bottleneck-shaped module sequentially into each frozen backbone layer. Most transformer-based REC models [3, 7, 18, 19] use pre-trained Vision Encoder and Language Encoder to separately extract image and text features, which are then integrated to form multi-modality features for reasoning. A straightforward approach to apply adapter-tuning for REC is to insert the adapters into the transformer encoder layers to enhance fine-tuning efficiency (see Figure 1 (b)). However, this introduces two significant challenges: (1) Updating inserted adapters still requires backpropagation of gradients through the large pre-trained vision and language models, placing a heavy burden on GPU memory during fine-tuning (see Figure 1 (b)). (2) Vision-language foundation models are pre-trained separately, each with its own structure and training data [20, 21]. Directly inserting vanilla adapters into them may lack cross-modality interaction in the shallow layers of the entire modal, resulting in sub-optimal vision-language alignment. This is particularly problematic for predicting referred objects in cases with complex semantic information, such as human actions and spatial relations (see Figure 4 "Without Interaction").
@@
To address these challenges, we propose a novel Multi-Modal Interactive Side-Tuning (M2IST) method that effectively strengthens vision-language alignment and enables parameter- and memory-efficient transfer to REC within the unified interactive side networks (see Figure 1 (c)). Specifically, we introduce Mixture of Multi-Modal Interactive Side-Adapters (M3ISAs), which incorporate Vision Expert Adapters (VEA), Language Expert Adapters (LEA), and Interaction Expert Adapters (IEA) into the side networks in parallel with the heavy encoders. VEA and LEA transfer pre-trained single-modality knowledge to the REC domain. IEA utilizes a linear layer for weight-sharing between image and text features, enabling progressive interaction between the referring sentence and input image. Such interaction aggregates channel-level vision-language alignment at shallow layers of the model, facilitating deep token-level vision-language alignment in deeper layers for improved performance. This elegant design achieves parameter-, memory-, and time-efficient intra- and inter-modality representation transfer for REC.
@@
We conduct extensive experiments on RefCOCO [22], RefCOCO+ [22], and RefCOCOg [23, 24] to demonstrate the effectiveness and efficiency of M2IST for REC. Experimental results show that M2IST achieves the optimal performance-parameter-memory trade-off compared to most full fine-tuning methods and other PETL methods. By applying our M2IST method, a standard transformer-based REC model can requiring only 2.11% of the tunable parameters, 39.61% of the GPU memory and 63.46% of the fine-tuning time compared to full fine-tuning, while still achieving competitive performance (see Figure1). With the sufficient vision-language interaction strengthened by our M3ISAs, our method can accurately locate the referred objects for various complex cases (see Figure 4).
To summarize, our main contributions are three-fold:
  1. 1.
    We propose M2IST, a novel Multi-Modal Interactive Side-Tuning method for referring expression comprehension (REC), which effectively addresses the challenges of insufficient multi-modal interaction and high GPU memory consumption in applying parameter-efficient transfer learning (PETL) to REC.
  2. 2.
    We design Mixture of Multi-Modal Interactive Side Adapters (M3ISAs), seamlessly integrating pre-trained vision and language encoders, enabling parameter-, memory-, and time-efficient tuning within a unified interactive side network for REC.
  3. 3.
    We conduct empirical studies on the application of PETL methods in REC, highlighting their limitations in practical scenarios. Extensive experiments on three widely-used benchmarks validate the effectiveness of M2IST, achieving the optimal trade-off between performance and efficiency compared to full fine-tuning and other PETL methods, with significantly reduced GPU memory usage and fine-tuning time.
II Related Work
II-A Referring Expression Comprehension
Referring expression comprehension (REC) [1, 2, 3, 25, 7, 26, 4, 9] aims to locate the specific objects in images based on textual descriptions.
Early REC methods [1, 27, 28] follow a two-stage pipeline that first uses a pre-trained object detector (e.g., Faster R-CNN [29]) to generate a set of sparse object proposals, which are then ranked by their similarity to the given textual description. MAttNet [1] uses Faster R-CNN to generate object proposals, then scores them based on the referring expression to find the most relevant object. RvG-Tree [30] builds a relational visual graph to capture object interactions and ranks proposals according to their relevance to the expression, improving grounding accuracy. However, these two-stage REC methods heavily rely on the quality of the object proposals and cannot directly predict the referred object region. One-stage anchor-based methods [2, 31, 32, 33] have been developed to eliminate the proposal generation step, directly predicting object bounding boxes from predefined anchors [34]. FAOA [2] utilizes the YOLOv3 detector [34], integrating it with an encoded language vector to ground the referred regions. MRLN [4] introduces three modules: feature-feature relational learning, feature-task relational learning, and task-task relational learning, to enhance the collaborative learning of REC, effectively reducing prediction inconsistency in multi-task learning. Recently, transformer-based methods [3, 35, 26, 7, 9] have demonstrated superior performance by implicitly modeling cross-modality relationships within a unified architecture. The pioneering work TransVG [3] applies a stack of transformer encoders to perform feature extraction and multi-modal fusion for REC. VGTR [35] employs a transformer encoder-decoder architecture to jointly reason over visual and textual inputs, grounding the referred object without relying on pre-trained detectors or word embeddings. Dynamic M-DETR [9] uses a dynamic multi-modal transformer decoder to adaptively sample visual features and perform text-guided decoding for REC.
As REC models continue to scale up, their performance has improved. However, this performance gain comes at the cost of increased computational cost, which demands larger GPU memory for more parameter fitting (see Figure 1 (a)).
II-B Parameter-efficient Transfer Learning
Parameter-efficient transfer learning (PETL) [12, 13, 14, 15] has emerged as a promising alternative to fully fine-tuning pre-trained models for downstream tasks. By updating only a minimal subset of parameters, PETL methods balance performance and computational efficiency.
Recent PETL methods can be classified into two categories. The first category is updating additional parameters in modules inserted into the model [12, 15, 36] or appended to the input data [14, 37, 38]. Adapter [12] incorporates a bottleneck module into each Transformer layer, positioned after both the Multi-Head Attention (MHA) and the Feed-Forward Networks (FFN). AdaptFormer [15] embeds the adapter module parallel to the FFN in each encoder of a Vision Transformer. VPT [14] appends learnable visual vectors to the input sequences (VPT-Shallow) or to the input of each transformer encoder layer (VPT-Deep). The second category involves decomposing weight matrices into two low-rank matrices and updating only the small factorization matrices [13, 39]. As a pioneering work, LoRA [13] integrates a tunable pair of low-rank decomposed weight matrices into each encoder layer of the pre-trained networks. FacT [39] incorporates tunable factorized weight matrices into each layer of the pre-trained networks. There is also growing interest in adapter-based PETL methods for vision-language tasks, including text-image retrieval [11] and text-video retrieval [40, 41]. Most of these methods aim to achieve effective cross-modality interaction while maintaining parameter efficiency.
However, existing PETL methods still face substantial GPU memory consumption during the fine-tuning stage, as gradients must propagate through the heavy pre-trained encoders for REC (see Figure 1 (b)).
II-C Memory-efficient Transfer Learning
Memory-efficient transfer learning (METL) [42, 43, 44] aims to reduce memory costs on GPUs during fine-tuning. Existing METL methods typically employ a side network for single-modality knowledge transfer, focusing on either NLP [43] or CV [44] downstream tasks. Side-Tuning [42] utilized an additional side network that adds its representation to the backbone network in the last layer. LST [43] adopted a separate and lightweight side network with the same architecture as the pre-trained model but reduced each layer dimension by a pre-defined reduction factor. DTL [44] disentangled the weights update from the pre-trained backbone network by proposing a lightweight side network, which achieved high accuracy in classification with low GPU memory usage.
In this work, inspired by the existing efforts, our M2IST bridges the pre-trained Vision Encoder and Language Encoder through the unified side interactive networks, enabling a parameter-, memory-, and time-efficient transfer to the REC task (see Figure 1 (c)).
@@ II-C Memory-efficient Transfer Learning
Figure 2: Overall architecture of M2IST. M2IST freezes the pre-trained Vision Encoder (blue branch) and Language Encoder (green branch), while updating M3ISAs on side networks (pink branch). M3ISAs comprise IEA for bridging the pre-trained dual encoders to enable cross-modality interactions, and VEA/LEA for transferring pre-trained single-modality representations to adapt to the REC domain. By avoiding backpropagation through the heavy encoders (red dashed arrow), M2IST enables parameter-, memory-, and time-efficient tuning for the task of referring expression comprehension.
III Methodology
In this section, we present our M2IST in detail. First, we briefly overview the base architecture for referring expression comprehension in Section III-A. Then, we elaborate on the designs of our efficient tuning method M2IST and its core component M3ISA in Section III-B. Finally, we provide an in-depth analysis of some advantages of M2IST in Section III-C.
III-A Base Architecture
@@
We apply a standard transformer-based REC model as our base architecture, shown in Figure 1 (a), which comprises: (1) a Vision Encoder, (2) a Language Encoder, and (3) a Vision-language Encoder.
Vision Encoder
@@
We adopt a DETR-based [20] encoder as our Vision Encoder, which comprises a ResNet [45] and a stack of transformer encoder layers to encode the image into high-quality vision embeddings. Specifically, given an input image $\bm{z}_{0}\in\mathbb{R}^{H_{0}\times W_{0}\times 3}$, the ResNet is utilized to generate a 2D feature map $\bm{z}\in\mathbb{R}^{H\times W\times C}$, where $H_{0}$ and $W_{0}$ denote the height and width of the input image, $H=\frac{H_{0}}{32}$, $W=\frac{W_{0}}{32}$, and $C=2048$ represents the channel dimension. Then, a $1\times 1$ convolutional layer reduces the $C$ to $C_{v}=256$, producing $\bm{z}^{\prime}\in\mathbb{R}^{H\times W\times C_{v}}$. We flatten the feature map $\bm{z}^{\prime}$ into a sequence of 1D vectors (i.e., vision tokens) $\bm{z}_{v}\in\mathbb{R}^{N_{v}\times C_{v}}$, where $N_{v}=H\times W$ indicates the number of tokens. Sequentially, these vision tokens added with positional encodings are fed into a stack of 6 transformer encoder layers, which then output the enhanced vision embeddings $\bm{f}_{v}\in\mathbb{R}^{N_{v}\times C_{v}}$ incorporating global context of the image.
Language Encoder
@@
We employ an off-the-shelf language model BERT [21], comprising a stack of transformer encoder layers, as our Language Encoder. Specifically, given the input text, each word ID is converted into a one-hot vector, which is then tokenized into a sequence of language tokens. These language tokens, concatenated with a [CLS] token at the beginning and a [SEP] token at the end, are input to 12 transformer encoder layers to sequentially model contextual relationships. Similar to the Vision Encoder, Language Encoder finally outputs the enhanced language embeddings $\bm{f}_{l}\in\mathbb{R}^{N_{l}\times C_{l}}$, where $N_{l}$ and $C_{l}=768$ represent the number and channel dimension of language tokens, respectively.
Vision-language Encoder
@@
We use a transformer-based encoder [46] as our Vision-language Encoder (V-L Encoder) to thoroughly fuse the multi-modality embeddings and predict the bounding box of the referred object. Specifically, the enhanced vision embeddings $\bm{f}_{v}\in\mathbb{R}^{N_{v}\times C_{v}}$ and language embeddings $\bm{f}_{l}\in\mathbb{R}^{N_{l}\times C_{l}}$ are first projected into the joint embeddings $\bm{f^{\prime}}_{v}\in\mathbb{R}^{N_{v}\times C_{p}}$ and $\bm{f^{\prime}}_{l}\in\mathbb{R}^{N_{l}\times C_{p}}$, sharing the same channel dimension $C_{p}=256$. The joint embeddings, along with a learnable [REG] token, are then fed into a stack of 6 transformer encoder layers to fuse the cross-modality embeddings and output the [REG] token. Finally, a prediction head, implemented as a Multi-layer Perceptron with two 256-dim hidden layers and a linear output layer, receives the [REG] token and regresses it to the 4-dim box coordinates for the referred object.
III-B Multi-Modal Interactive Side-Tuning: M2IST